OpenStudy (thadds2003):

Please help, @pooja195.

pooja195 (pooja195):

with?

pooja195 (pooja195):

What expression?

OpenStudy (thadds2003):

the fraction.

pooja195 (pooja195):

What fraction?

pooja195 (pooja195):

oh nvm i see it you need to find the LCD

OpenStudy (thadds2003):

ok, getting it...... 210? @pooja195

pooja195 (pooja195):

No... What is the lowest number that 15 and 6 both multiply into

OpenStudy (thadds2003):

90.

pooja195 (pooja195):

How did you get that?

OpenStudy (thadds2003):

its easy, just multiply 15 *6 and get the answer, 90.

OpenStudy (thadds2003):

@pooja195

OpenStudy (thadds2003):

i got it, pooja. next step?

pooja195 (pooja195):

We dont multiply them

OpenStudy (thadds2003):

?

pooja195 (pooja195):

We need to find the LEAST common denominator.

OpenStudy (thadds2003):

wich is 90....... or no?

pooja195 (pooja195):

90/2=?

OpenStudy (thadds2003):

4.5, @pooja195

pooja195 (pooja195):

*90/3

OpenStudy (thadds2003):

oh, made a mistake? well thew answer to that is 30.

OpenStudy (thadds2003):

* the

pooja195 (pooja195):

\[\huge~\rm~\frac{ 14(2) }{ 15(2) }-\frac{ 1(5) }{ 6(5) }\rightarrow \frac{ 28-5 }{ 30 }\]

OpenStudy (thadds2003):

23/30 = 0.766666667. had to use the calc, couldent solve this one. :/

pooja195 (pooja195):

Leave it as a fraction since thats how your questions wants it